Five less than six times a number is 61
1
Expert's answer
2020-02-10T10:42:01-0500

Solution:

Let the number=n

So, five less than six times a number is 61:

6n-5=61

6n=61+5

6n=66

n=66/6

n=11

Answer: 11.